Use what you need of the following set of information to answer the question:

S0 = $1.1420/€            S0 = £0.7776/$                        S0 = $1.2860/£       

                        E (S1 (3-months)) = $1.1280/€                              F3-months = $1.1590/€

                 Where: € = euro; £ = British pound; E ( ) = Expected.

  1. If you have $248,000,000 how many British pounds would you be able to buy for immediate delivery given the above exchange rates? (Assume no bid/ask spread)
  2. If you have €640,000,000 how many dollars would you be able to buy for immediate delivery given the above exchange rates? (Assume no bid/ask spread)
  3. If you have $320,000,000 how many euros would you be able to buy for 3-month delivery given the above exchange rates? (Assume no bid/ask spread)

(Show all your calculations and explain your work. All currency amount answers should be presented with two decimal places, all currency rates with the same number of decimals they are given, and all rates of change and premiums/discounts with four decimal places)
Answer #1
1) If you have $248,000,000 how many British pounds would you be able to buy for immediate delivery given the above exchange rates? (Assume no bid/ask spread)
The spot rate for GBP should be used. The # of GBP that can be purchased = 248000000/1.2860 = £        19,28,46,034.21
2) If you have €640,000,000 how many dollars would you be able to buy for immediate delivery given the above exchange rates? (Assume no bid/ask spread)
The spot rate for Euro should be used. The # of Euro that can be purchased = 640000000/1.1420 = €        56,04,20,315.24
3) If you have $320,000,000 how many euros would you be able to buy for 3-month delivery given the above exchange rates? (Assume no bid/ask spread)
The forward rate for Euros should be used. The # of Euros
that can be bought for delivery after 3 months = 320000000/1.1590 = €        27,61,00,086.28
It is assumed at F3 represents the 3 month forward rate for Euro.